OpenStudy (anonymous):

Please help! Michael makes 90% of free throws. if he shoots 20 what is the probability he makes... all 20? exactly 18? at least 18?. his chance of making each one is independent of the other shots

OpenStudy (anonymous):

Since he makes 90% of throws, 90% of 20 = 90/100 * 20 = 18. So P(getting exactly 18) = 1 <sure event> Also P(getting at least 18) = 1 <sure event> P(getting all 20) = 18/20 = 9/10

OpenStudy (anonymous):

yes, but since he makes only 90% of his free throws, the chance of him making 100% of them is 0 right?

OpenStudy (anonymous):

I think you're right. So P(getting all 20) = 0 since it is an impossible event.

OpenStudy (anonymous):

okay :) thankss

OpenStudy (anonymous):

The answer you got to this was not correct. If he has a 90% chance of making each free-throw then there is indeed a non-zero probability he'll make all 20.

OpenStudy (anonymous):

Specifically it's .9^20, or a 12.2% chance.